# Content-based Wake-up for Energy-efficient and Timely Top-$k$ IoT Sensing Data Retrieval

### v1

@@ III-B Baseline 1: RR-scheduling
As a main baseline, we apply RR-scheduling, in which each node transmits its measurement according to a Time Division Multiple Access (TDMA)-like policy. In this scheme, the sink transmits a single wake-up signal to trigger all nodes equipped with the wake-up receiver at time $t_{\mathrm{sch}}$, which is $NL$ slots prior to the deadline. The nodes receiving the wake-up signal then sample and transmit their measurements in order, i.e., node $j=0,1,\ldots,N-1$, samples physical process and transmits its measurement at $t_{\mathrm{sch}}+jL$. The $k$-QAoI of RR-scheduling can be expressed as:
@@ III-C Baseline 2: Probability $q$-based Random Wake-up
As a second baseline, we consider the probability $q$-based random wake-up, which we call $q$-Wu. In this scheme, the sink transmits a wake-up signal $\zeta$ slots prior to the deadline. Each sensor node receiving the wake-up signal randomly decides with probability $q$ whether it wakes up (regardless of its content), and otherwise stays silent. The node woken up transmits its data following the $p$-persistent CSMA. The probability that $\nu$ out of $N$ nodes wake up can be expressed as $P_{x}(\nu)=\binom{N}{\nu}q^{\nu}(1-q)^{N-\nu}$. By applying Eq. (11), the probability $\nu_{s}$ out of $\nu$ nodes succeeds in data transmission by the deadline $T$, given $\nu$ and $\zeta$, can be described as $P_{s}(\nu_{s}|\nu,\zeta)$. Finally, the probability that $c$ out of $\nu_{s}$ nodes belonging to the top-$k$ set at $t_{s}$ succeeds in data transmission can be expressed as
@@ III-D Baseline 3: Genie-aided Scheme
$\tilde{\Delta}_{k}(k)=\frac{1}{k}\sum_{w=1}^{k}\min(f_{s}(wL),A_{\mathrm{max}}).$ (21)
@@
Accordingly, the lower bound of the energy consumption can be derived as follows. Let $O(K_{1})$ be the total energy consumption when the sink collects data from $K_{1}\geq k$ sensor nodes under the same access protocol. Then, the following statement is always true: $O(K_{1})\geq O(k)$. Therefore, Eq. (22) is the lower bound of total energy consumption under the lower bound of $k$-QAoI, in which the nodes in $\Omega_{k}$ activate their main radio only in their allocated slots.
@@ IV-A Problem Formulation
As obtaining the exact optimal values is computationally difficult, we rely on the heuristic solution, as we will describe in Sec. IV-B, in which we obtain the set of optimal parameters with two steps based on grid search.
IV-B Heuristic Solution
Step 1. Optimizing the transmission probability $p$
As an optimal parameter of $p$ depends on the number of wake-up nodes, we first derive the optimal parameters of $p$ heuristically, based on the following equations of data collection delay derived in [41, 5]:
$T_{d}(w)=\sum_{m=1}^{w}\frac{L-(L-1)(1-p)^{m}}{(1-e_{c})mp(1-p)^{m-1}}{\delta}.$ (28)
@@ IV-B Heuristic Solution
As a special case, this work assumes $T_{d}(0)=0$. Here, we optimize the parameter of $p$, denoted as $p_{\mathrm{opt}}$ in terms of data collection delay. This is because a smaller delay is important to realize information freshness at the deadline, by which the sink can activate nodes at a much later time to obtain smaller $k$-QAoI. Specifically, we tune the value of $p$ in accordance with the number of wake-up nodes $w$, so that the expected data collection delay can be minimized, as shown below.
$\ p_{\mathrm{opt}}(w)=\min_{p\in(0,1]}T_{d}(w).$ (29)
Here, we obtain the optimal value of $p$ in a heuristic way using Eq. (28), in which we change the value of $p$ out of the range from 0.001 to 1 with a step of 0.001.
Step 2. Optimizing the parameters $\zeta$ and $V_{\mathrm{th}}$ for CoWu signal
@@
We then obtain the optimal set of parameter $\mathbb{C}(N)=\{\zeta,V_{\mathrm{th}}\}$ when we apply $p_{\mathrm{opt}}(w)$ shown in Eq. (29) based on grid search, which are required to satisfy
@@
where we select the value of $V_{\mathrm{th}}$ out of the range from 0 to 50 with a step of 0.5, and that of $\zeta$ out of the range from 10 to 1000 with a step of 10.
@@ IV-C Maximum $k$
We resort to a heuristic approach as in Sec. IV-B to obtain the maximum value of $k$, in which results of average $k$-QAoI and total energy consumption are evaluated for a wide range of parameters $\{\zeta,V_{\mathrm{th}}\}$. Here, we apply $p_{\mathrm{opt}}(w)$ in Eq. (29) and select the value of $V_{\mathrm{th}}$ out of the range from 0 to 50 with a step of 2, and that of $\zeta$ out of the range from 50 to 500 with a step of 50 for the different number of nodes $N$.
@@ V-A Achievable Region for Different Wake-up Scheme
[…]
@@ V-B1 Impact of Timing of CoWu Signal $\zeta$
First, we investigate the impact of the timing of CoWu signal in terms of $k$-QAoI. Figs. 3 and 4 show $k$-QAoI of CoWu against $\zeta$ for the linear age function and for the exponential age function defined in Eq. (4) with $\alpha=0.02$, respectively, where we set $N=100$, $k=5$, $e_{c}=0$, $p=0.0606$, $\Gamma=1000$, and $V_{\mathrm{th}}~{}=~{}46,~{}48$. From these figures, we can first see that the results for CoWu obtained with theoretical analysis coincide with simulation results for both the linear age case and exponential age case, which validates our analysis. Next, from Figs. 3 and 4, we can also see that there is an optimal value of $\zeta$ in terms of $k$-QAoI when the value of $V_{\mathrm{th}}$ is fixed. Let us denote the optimal value of $\zeta$ as $\zeta_{\mathrm{opt}}$. For $\zeta<\zeta_{\mathrm{opt}}$, we can see that the $k$-QAoI becomes larger as $\zeta$ decreases, because the number of nodes that fail their data transmission by the deadline $T$ increases due to the congestion. For $\zeta>\zeta_{\mathrm{opt}}$, most users complete their transmission, but the sensed values become obsolete at the deadline, also leading to an increase of $k$-QAoI. This result illustrates the importance of the timing of the wake-up signal so as to minimize the $k$-QAoI at the deadline. Further, from Fig. 4, we can see that the $k$-QAoI of the exponential case is much larger than that of the linear age case shown in Fig. 3 and it reaches maximum age $A_{\max}$ when the value of $\zeta$ is equal to or larger than 450. This is because the collected data in the exponential age case becomes obsolete much faster than the linear age case (c.f. Sec. V-B2). Finally, we can clearly see CoWu achieves smaller $k$-QAoI than RR-scheduling by transmitting a wake-up signal at the right timing against the deadline, i.e., by selecting the appropriate parameter of $\zeta$.
@@ V-B2 Timing Analysis Considering Different Characteristics of Physical Process
Figs. 5 and 6 show the optimal timing of CoWu signal and optimal $k$-QAoI against the value of $\alpha$ for the exponential age in Eq. (4), where we set $N=100$, $k=5$, $e_{c}=0$, $p=0.0606$, $\Gamma=1000$, and $V_{\mathrm{th}}=46$. Here, the optimal timing of wake-up signaling is obtained based on the heuristic approach. Note that the value of $\alpha$ represents the different speeds of changes of the physical process: smaller (larger) $\alpha$ represents a slowly (fast) changing physical process. From Fig. 5, we can see that, in the range of large $\alpha$, the value of $\zeta_{\mathrm{opt}}$ becomes smaller, in which each node should transmit data later time in order to prevent data staleness at the deadline. This is because earlier transmission of wake-up signal leads to data staleness when the speed of change of the observed physical process is higher. On the other hand, when $\alpha$ is smaller, the earlier transmission of CoWu signal (a larger value of $\zeta$) is the best strategy in order to reduce $k$-QAoI, by which the probability of succeeding in data transmission by the deadline for each wake-up node increases. Here, with a smaller range of $\alpha$, e.g. ($\alpha<0.01$), we can see that the $\zeta_{\mathrm{opt}}$ becomes smaller as the value of $\alpha$ decreases. In this range, thanks to the slowly changing physical process, failure of data transmission does not significantly deteriorate $k$-QAoI as $f_{s}(\Gamma)$ is small. Thus, the sink can transmit a wake-up signal at a later time to collect fresher information from the sensor nodes to reduce $k$-QAoI. The result of Fig. 5 brings us insight into the importance of the timing of wake-up signaling in accordance with the characteristics of physical process observed by each sensor.
@@ V-C1 Investigation of the Impact of Packet Error
Fig. 8 shows $k$-QAoI of CoWu and RR-scheduling against the value of $e_{c}$, where we set $N=100$, $k=10$, $V_{\mathrm{th}}=46$, $\zeta=150$, and the optimal transmission probability of $p$ obtained by Eq. (29), with linear age scenario in (4), and evaluate the cases of $\Gamma=1000$ and $\Gamma=5000$. First, from Fig. 8, we can see that $k$-QAoI of each scheme becomes larger as the value of $e_{c}$ increases. In CoWu, the node that fails to transmit data due to the single packet error continues to attempt data transmission following $p$-persistent CSMA, which increases the contention time and decreases the probability of succeeding in data transmission by the deadline, which is why $k$-QAoI increases as the value of $e_{c}$ increases. On the other hand, in RR-scheduling, if a single packet error occurs for a time slot, the sink can not retrieve the information of the observation values of this failed packet, by which $k$-QAoI of RR-scheduling becomes larger. Next, focusing on the case where the penalty for the data transmission failure is relatively small, i.e., the case for $\Gamma=1000$, we can see that CoWu achieves smaller $k$-QAoI than that of RR-scheduling. This is because, in CoWu, each node can retransmit data following the operation of $p$-persistent CSMA when experiencing a single packet error, which increases the chance of success in data transmission. On the other hand, in RR-scheduling, as mentioned above, the $k$-QAoI becomes higher by the penalty of $\Gamma$ when the single packet error occurs. Next, focusing on the case for $\Gamma=5000$, first, we can see that the slope of $k$-QAoI of both CoWu and RR-scheduling is larger than that of the case for $\Gamma=1000$. This is simply because of the increasing penalty value of $\Gamma$ for the data collection failure. We can also see that $k$-QAoI of CoWu becomes larger than that of RR-scheduling for the case of $e_{c}=0$ and $\Gamma=5000$. When the value of the $\Gamma$ is higher, it is important for the sink to collect the fresh data from the sensor nodes. In RR-scheduling, all nodes can reliably transmit data to the sink when $e_{c}=0$ because there is no collision, while in CoWu, data transmission failure may occur due to the congestion or inappropriate threshold of CoWu, by which the $k$-QAoI becomes larger, especially when the value of $\Gamma$ is larger. That is why $k$-QAoI of CoWu becomes larger than that of RR-scheduling when $\Gamma=5000$ and $e_{c}=0$. However, even if the value of $\Gamma$ is higher, such as $\Gamma=5000$, we can see that the CoWu can achieve much smaller $k$-QAoI than that of RR-scheduling when the value of $e_{c}$ is larger.
@@
Figure 8: $k$-QAoI of CoWu and RR-scheduling against $e_{c}$ (Linear age).

### v2

@@ III-B Baseline 1: RR-scheduling
As a main baseline, we apply RR-scheduling, in which each node transmits its measurement according to a Time Division Multiple Access (TDMA)-like policy. In this scheme, the sink transmits a single wake-up signal to trigger all nodes equipped with the wake-up receiver at time $t_{\mathrm{sch}}$, which is $NL$ slots prior to the deadline. The nodes receiving the wake-up signal then sample and transmit their measurements in order, i.e., node $j=0,1,\ldots,N-1$, samples physical process and transmits its measurement at $t_{\mathrm{sch}}+jL$. To derive the $k$-QAoI of RR-scheduling, we first calculate the expected QAoI of nodes observing $i\in[1,k]$-th largest values. The transmission time and sampling time of the nodes observing $i$-th largest values is a random variable, uniformly distributed between the duration of [$L$, $NL$] with an interval of $L$. Based on this observation, the expected QAoI of $i$-th largest node can be expressed as $\frac{1}{N}\sum_{w=1}^{N}(1-e_{c})\min(f_{s}(wL),A_{\mathrm{max}})+e_{c}\min(f_{s}(\Gamma),A_{\mathrm{max}})$. Finally, based on the definition of Eq. (6), the expected $k$-QAoI of RR-scheduling can be expressed as in Eq. (17), in which the value of $k$ is canceled between denominator and numerator:
@@ III-C Baseline 2: Probability $q$-based Random Wake-up
As a second baseline, we consider the probability $q$-based random wake-up, which we call $q$-Wu. In this scheme, the sink transmits a wake-up signal $\zeta$ slots prior to the deadline. Each sensor node receiving the wake-up signal randomly decides with probability $q$ whether it wakes up (regardless of its content), and otherwise stays silent. The node woken up transmits its data following the $p$-persistent CSMA. The probability that $\nu$ out of $N$ nodes wake up can be expressed as $P_{x}(\nu)=\binom{N}{\nu}q^{\nu}(1-q)^{N-\nu}$. By applying Eq. (12), the probability $\nu_{s}$ out of $\nu$ nodes succeeds in data transmission by the deadline $T$, given $\nu$ and $\zeta$, can be described as $P_{s}(\nu_{s}|\nu,\zeta)$. Finally, the probability that $c$ out of $\nu_{s}$ nodes belonging to the top-$k$ set at $t_{s}$ successfully transmit can be expressed as
@@ III-D Baseline 3: Genie-aided Scheme
$\begin{split}\tilde{\Delta}_{k}(k)&=\frac{1}{k}\sum_{w=1}^{k}(1-e_{c})\min(f_{s}(wL),A_{\mathrm{max}})\\ &+e_{c}\min(f_{s}(\Gamma),A_{\mathrm{max}}).\end{split}$ (22)
@@
Accordingly, the lower bound on the energy consumption can be derived as follows. Here, we are interested in a lower bound on the energy consumption given $k$ that achieves the lower bound of $k$-QAoI $\tilde{\Delta}_{k}(k)$ in Eq. (22). Let $O(K_{1})$ be the total energy consumption when the sink collects data from $K_{1}\geq k$ sensor nodes under the same access protocol. Then, the following statement is always true: $O(K_{1})\geq O(k)$. Therefore, Eq. (23) is the lower bound of total energy consumption under the lower bound of $k$-QAoI, in which the nodes in $\Omega_{k}$ activate their main radio only in their allocated slots.
Finally, it is worth emphasizing that the genie-aided scheme cannot be implemented in practice, since it relies on the assumption of having a perfect and instantaneous observation of the status of the nodes, including those that belong to the top-$k$ set.
@@ IV-A Problem Formulation
As obtaining the exact optimal values is computationally and analytically difficult, we introduce an approximate solution, as we will describe in Sec. IV-B, in which we obtain the set of optimal parameters with two steps based on an analysis and a grid search.
IV-B Approximate Solution
In order to tackle the problem mentioned in Sec. IV-A, we introduce an approximate solution to derive the optimal sets of parameter $\{V_{\mathrm{th}},p,\zeta\}$ for Eqs. (24)–(28). Specifically, we first obtain the optimal transmission probability $p$ analytically (c.f. Sec. IV-B1) and then using the optimal transmission probability, we obtain the optimal parameter of $\{\zeta,V_{\mathrm{th}}\}$ with grid-search (c.f. Sec. IV-B2).
IV-B1 Step 1. Optimizing the Transmission Probability $p$
@@ IV-B Approximate Solution
Here, we first optimize the parameter of $p$ in terms of data collection delay. This is because a smaller delay is important to realize information freshness at the deadline, by which the sink can activate nodes at a much later time to obtain smaller $k$-QAoI.
In the OSD model, the number of active users gradually decreases as packets are successfully transmitted. In [42], in order to analyze the delay and energy consumption of the nodes operating with $p$-persistent CSMA protocol under OSD, the authors introduced the notion of an “epoch,” which is defined as the time required for a single sensor node succeeding in data transmission when the $m$ active nodes compete for the channel. Then, the authors analyzed the expected time for an epoch with $m$ active nodes analytically, denoted as $\mathbb{E}[T_{m}]$, which can be expressed as [42, 5]:
$\mathbb{E}[T_{m}]=\frac{L-(L-1)(1-p)^{m}}{(1-e_{c})mp(1-p)^{m-1}}{\delta}.$ (29)
Then, data collection delay when we employ $p$-persistent CSMA in OSD can be expressed as follows [42, 5]:
$T_{d}(w)=\sum_{m=1}^{w}\mathbb{E}[T_{m}],$ (30)
where this work assumes $T_{d}(0)=0$, as a special case.
From Eq. (30), we can observe that, to minimize the data collection delay $T_{d}(w)$ in Eq. (30), we must minimize the delay $\mathbb{E}[T_{m}]$ in Eq. (29). As $\mathbb{E}[T_{m}]$ in Eq. (29) depends on the value of $p$ and $m$, we take an approach to optimize $p$ in $\mathbb{E}[T_{m}]$ for each $m$ step-by-step so that $\mathbb{E}[T_{m}]$ can be minimized. Let us denote the optimal transmission probability $p$, which minimizes $\mathbb{E}[T_{m}]$ as $p_{\mathrm{opt}}^{*}(m)$. Then, we can formulate the optimization problem for each $m\in\{1,2,\ldots,w\}$ as follows:
$p_{\mathrm{opt}}^{*}(m)=\mathop{\rm arg~{}min}\limits_{p\in(0,1]}\mathbb{E}[T_{m}].$ (31)
We analytically solve the problem in Eq. (31) to obtain optimal transmission probability. Specifically, we set the first order derivative of $\mathbb{E}[T_{m}]$ with respect to $p$ to be 0, i.e., $\frac{\mathrm{d}\mathbb{E}[T_{m}]}{\mathrm{dp}}=0$, for the case of $L>1$. After several steps of the calculation, we can obtain the following results: $(1-p)^{m}=\frac{L}{L-1}(1-mp)$. Then, by applying the second order approximation of Taylor expansion around $p=0$, we can obtain the following approximation: $(1-p)^{m}\approx 1-mp+\frac{m(m-1)}{2}p^{2}$. Finally, we can obtain the approximate optimal transmission probability for the case of the active number of wake-up nodes $m>1$, as in [42], as $\frac{\sqrt{m^{2}+2m(m-1)(L-1)}-m}{m(m-1)(L-1)}$. When $m=1$, with Eq. (30), data collection delay can be expressed as $\mathrm{E}[T_{1}]=\frac{L-1}{1-e_{c}}\delta+\frac{1}{p(1-e_{c})}\delta$. Clearly, this is a decreasing function for $p\in(0,1]$, which can be minimized when $p=1$. By combining these results, we can express the optimal transmission probability for each $m$ as follows:
$p_{\mathrm{opt}}^{*}(m)=\begin{cases}1,&m=1,\\ \approx\frac{\sqrt{m^{2}+2m(m-1)(L-1)}-m}{m(m-1)(L-1)},&m>1.\end{cases}$ (32)
The $k$-QAoI and total energy consumption of CoWu applying dynamic transmission probability $p_{\mathrm{opt}}^{*}(m)$ can be calculated by substituting $p_{\mathrm{opt}}^{*}(m)$ into Eqs. (9) and (10), and into the Eq. (15) for each $m$.
Note that, in order to apply the optimal transmission probability $p_{\mathrm{opt}}^{*}(m)$ in Eq. (32), each node must know the number of active nodes given time. In practice, this can be realized through the reception of broadcasted ACK. Specifically, given the number of wake-up nodes $w$, each node updates the number of active nodes and its transmission probability upon the reception of ACK.
IV-B2 Step 2. Optimizing the parameters $\zeta$ and $V_{\mathrm{th}}$ for CoWu signal
@@
We then obtain the optimal set of parameter $\mathbb{C}(N)=\{\zeta,V_{\mathrm{th}}\}$ when we apply $p_{\mathrm{opt}}^{*}$ shown in Eq. (32) based on grid search, which are required to satisfy
@@
where we select the value of $V_{\mathrm{th}}$ out of the range from 0 to 50 with a step of 0.5, and that of $\zeta$ out of the range from 10 to 1000 with a step of 10. Note that we rely on grid-search methods because the problem in Eq. (33) is non-convex and cannot be solved analytically.
@@ IV-C Maximum $k$
Because Eqs. (34)–(36) are non-convex, we resort to the approximate solution as in Sec. IV-B to obtain the maximum value of $k$, in which results of average $k$-QAoI and total energy consumption are evaluated for a wide range of parameters $\{\zeta,V_{\mathrm{th}}\}$. Here, we apply the optimal dynamic transmission probability $p_{\mathrm{opt}}^{*}$ obtained in Eq. (32), and select the value of $V_{\mathrm{th}}$ out of the range from 0 to 50 with a step of 2, and that of $\zeta$ out of the range from 50 to 500 with a step of 50 for the different number of nodes $N$.
@@ V-A Achievable Region for Different Wake-up Scheme
$0001{,}$  [^0] Total energy consumption [mJ]k-QAoICoWu (Theory, $\zeta=250$)CoWu (Simulation, $\zeta=250$)$q$-Wu (Theory, $\zeta=250$)$q$-Wu (Simulation, $\zeta=250$)Round-robin schedulingGenie-aided Scheme
@@ V-B1 Impact of Timing of CoWu Signal $\zeta$
First, we investigate the impact of the timing of CoWu signal in terms of $k$-QAoI. Figs. 4 and 5 show $k$-QAoI of CoWu against $\zeta$ for the linear age function and for the exponential age function defined in Eq. (4) with $\alpha=0.02$, respectively, where we set $N=100$, $k=5$, $e_{c}=0$, $p=0.0606$, $\Gamma=1000$, and $V_{\mathrm{th}}~{}=~{}46,~{}48$. From these figures, we can first see that the results for CoWu obtained with theoretical analysis coincide with simulation results for both the linear age case and exponential age case, which validates our analysis. Next, from Figs. 4 and 5, we can also see that there is an optimal value of $\zeta$ in terms of $k$-QAoI when the value of $V_{\mathrm{th}}$ is fixed. Let us denote the optimal value of $\zeta$ as $\zeta_{\mathrm{opt}}$. For $\zeta<\zeta_{\mathrm{opt}}$, we can see that the $k$-QAoI becomes larger as $\zeta$ decreases, because the number of nodes that fail their data transmission by the deadline $T$ increases due to the congestion. For $\zeta>\zeta_{\mathrm{opt}}$, most users complete their transmission, but the sensed values become obsolete at the deadline, also leading to an increase of $k$-QAoI. This result illustrates the importance of the timing of the wake-up signal so as to minimize the $k$-QAoI at the deadline. Further, from Fig. 5, we can see that the $k$-QAoI of the exponential case is much larger than that of the linear age case shown in Fig. 4 and it reaches maximum age $A_{\max}$ when the value of $\zeta$ is equal to or larger than 450. This is because the collected data in the exponential age case becomes obsolete much faster than the linear age case (c.f. Sec. V-B2). Finally, we can clearly see CoWu achieves smaller $k$-QAoI than RR-scheduling by transmitting a wake-up signal at the right timing against the deadline, i.e., by selecting the appropriate parameter of $\zeta$.
@@ V-B2 Timing Analysis Considering Different Characteristics of Physical Process
Figs. 6 and 7 show the optimal timing of CoWu signal and optimal $k$-QAoI against the value of $\alpha$ for the exponential age in Eq. (4), where we set $N=100$, $k=5$, $e_{c}=0$, $p=0.0606$, $\Gamma=1000$, and $V_{\mathrm{th}}=46$. Here, the optimal timing of wake-up signaling is obtained based on the grid search. Note that the value of $\alpha$ represents the different speeds of changes of the physical process: smaller (larger) $\alpha$ represents a slowly (fast) changing physical process. From Fig. 6, we can see that, in the range of large $\alpha$, the value of $\zeta_{\mathrm{opt}}$ becomes smaller, in which each node should transmit data later time in order to prevent data staleness at the deadline. This is because earlier transmission of wake-up signal leads to data staleness when the speed of change of the observed physical process is higher. On the other hand, when $\alpha$ is smaller, the earlier transmission of CoWu signal (a larger value of $\zeta$) is the best strategy in order to reduce $k$-QAoI, by which the probability of succeeding in data transmission by the deadline for each wake-up node increases. Here, with a smaller range of $\alpha$, e.g. ($\alpha<0.01$), we can see that the $\zeta_{\mathrm{opt}}$ becomes smaller as the value of $\alpha$ decreases. In this range, thanks to the slowly changing physical process, failure of data transmission does not significantly deteriorate $k$-QAoI as $f_{s}(\Gamma)$ is small. Thus, the sink can transmit a wake-up signal at a later time to collect fresher information from the sensor nodes to reduce $k$-QAoI. The result of Fig. 6 brings us insight into the importance of the timing of wake-up signaling in accordance with the characteristics of physical process observed by each sensor.
@@ V-C1 Investigation of the Impact of Packet Error
Fig. 9 shows $k$-QAoI of CoWu and RR-scheduling against the value of $e_{c}$, where we set $N=100$, $k=5$, $V_{\mathrm{th}}=46$, $\zeta=150$, and the optimal transmission probability of $p_{\mathrm{opt}}^{*}$ obtained by Eq. (32), with linear age scenario in Eq. (4), and evaluate the cases of $\Gamma=1000$ and $\Gamma=5000$.
First, from Fig. 9, we can see that the results for CoWu obtained with theoretical analysis coincide with the simulation results, which confirm the validity of our analysis applying optimal transmission probability. Next, from Fig. 9, we can see that $k$-QAoI of each scheme becomes larger as the value of $e_{c}$ increases. In CoWu, the node that fails to transmit data due to the single packet error continues to attempt data transmission following $p$-persistent CSMA, which increases the contention time and decreases the probability of succeeding in data transmission by the deadline, which is why $k$-QAoI increases as the value of $e_{c}$ increases. On the other hand, in RR-scheduling, if a single packet error occurs for a time slot, the sink can not retrieve the information of the observation values of this failed packet, by which $k$-QAoI of RR-scheduling becomes larger. Next, focusing on the case where the penalty for the data transmission failure is relatively small, i.e., the case for $\Gamma=1000$ in Fig. 9, we can see that CoWu achieves smaller $k$-QAoI than that of RR-scheduling. This is because, in CoWu, each node can retransmit data following the operation of $p$-persistent CSMA when experiencing a single packet error, which increases the chance of success in data transmission. On the other hand, in RR-scheduling, as mentioned above, the $k$-QAoI becomes higher by the penalty of $\Gamma$ when the single packet error occurs. Next, focusing on the case for $\Gamma=5000$ shown in Fig. 9, first, we can see that the slope of $k$-QAoI of both CoWu and RR-scheduling is larger than that of the case for $\Gamma=1000$ shown in Fig. 9. This is simply because of the increasing penalty value of $\Gamma$ for the data collection failure. When the value of $\Gamma$ is higher, it is important for the sink to collect the fresh data from the sensor nodes. In CoWu, data transmission failure may occur due to the congestion or inappropriate threshold of CoWu, by which the $k$-QAoI becomes larger, especially when the value of $\Gamma$ is larger. However, even if the value of $\Gamma$ is higher, such as $\Gamma=5000$, we can see that the CoWu can achieve much smaller $k$-QAoI than that of RR-scheduling especially when the value of $e_{c}$ is larger.
@@
Figure 9: $k$-QAoI of CoWu and RR-scheduling against $e_{c}$ (Linear age), where we apply $\Gamma=1,000$ and $\Gamma=5,000$.